# SDP bounds on quantum codes

### v1

@@ 3.5. Lovász theta number of a graph
3.6. State polynomial optimization
Given a state $\varrho$ and a set of observables $\{X_{i}\}^{m}_{i=1}$, denote by ${\langle X_{i}\rangle}_{\varrho}=\operatorname{tr}(X_{i}\varrho)$ their expectation values on $\varrho$. Consider now polynomials in products of operators and their expectations e.g., ${\langle(X_{1})^{4}\rangle}_{\varrho}\,X_{2}\,+\,(X_{3})^{2}$. We term these state polynomials in the variables $X_{i}$.
An optimization problem over the set of quantum states $\varrho$ and operators $X_{i}$ on a Hilbert space $\mathcal{H}$ with an objective state polynomial function $f$ and constraints $g_{j}$ then reads
(25) $\alpha^{*}=\max_{\mathcal{H},\{X_{i}\},\varrho}$ $\langle\,f(X_{1},\dots,X_{m})\,\rangle_{\varrho}$
subject to $g_{j}(X_{1},\dots,X_{m})\succeq 0\quad\text{for}\quad j=0,\dots,r\,.$
Note that Eq. (25) requires an optimization over the Hilbert space $\mathcal{H}$ and all operators and quantum states with support on $\mathcal{H}$ satisfying the inequality constraints. The framework of state polynomial optimization [33] (also known as scalar extension [53, 54]) provides a hierarchy of semidefinite programming approximations to Eq. (25) that converge.
To see how this works, consider a set of (non-commutative) letters $\{x_{i}\}^{m}_{i=1}$ and form words $w=x_{j_{1}}\dots x_{j_{m}}$. We associate to each word $w$ a scalar with the symbol ${\langle w\rangle}={\langle x_{j_{1}}\dots x_{j_{k}}\rangle}$. We refer to ${\langle w\rangle}$ as the “expectation value” of a word. These satisfy $w{\langle v\rangle}={\langle v\rangle}w$ and $({\langle v\rangle}w)^{*}={\langle v^{*}\rangle}w^{*}$ for all words $v,w$. Furthermore, denote the empty word as $1$, satisfying $1w=w1=w$ for all words $w$. The degree of a word is defined as the number of letters in it. The involution of a word is given by $(x_{1}\cdots x_{m})^{*}=(x^{*}_{m}\cdots x^{*}_{1})$. Products of words and expectation values of words are known as state monomials, e.g. $w_{i_{1}}{\langle w_{i_{2}}\rangle}\dots\langle w_{i_{m}}\rangle$. The degree of a state monomial is the sum over the degree of all words that compose it, including the words within expectation values. State polynomials are formed by a linear combination of monomials. The degree of a state polynomial is the largest degree of the monomials that it is composed of.
Naturally, state polynomials behave in the same manner as operators. A word $x_{i}$ and its expectation value ${\langle x_{i}\rangle}$ correspond to an operator $X_{i}$ and its expectation value ${\langle X_{i}\rangle}_{\varrho}$ from a state $\varrho$. In particular, the empty word $\mathds{1}$ corresponds to the identity, since $\mathds{1}W_{i}=W_{i}\mathds{1}=W_{i}$. Similarly, a state polynomial like $w={\langle x_{i}\rangle}x_{j}$ corresponds to $W={\langle X_{i}\rangle}_{\varrho}X_{j}$. However, a set of values ${\langle x_{i}\rangle}$ does not necessarily have to be compatible with a quantum state, i.e., it can be that there does not exist a $\varrho$ such that ${\langle x_{i}\rangle}\neq{\langle x_{i}\rangle}_{\varrho}$.
Let $\vec{w}^{(\ell)}$ be a vector that contains all state monomials up to degree $\ell$. Then, we can define the matrix $M^{(\ell)}_{ij}={\langle{w_{i}^{*}}^{(\ell)}w_{j}^{(\ell)}\rangle}$ with $M^{(\ell)}_{00}=1$. It is important to remark that if $M^{(\ell)}_{ij}$ comes from a state, then it can be written as
(26) $M^{(\ell)}_{ij}=\operatorname{tr}({W_{i}^{\dagger}}^{(\ell)}W_{j}^{(\ell)}\varrho)\,.$
Then, $M^{(\ell)}$ is a positive semidefinite matrix with $M^{(\ell)}_{00}=1$. This can be seen by writing the spectral decomposition of the state $\varrho=\sum_{v}p_{v}|\phi_{v}\rangle\langle\phi_{v}|$ with $p_{v}\geq 0$; then for each $\mathinner{|\phi_{v}\rangle}$ form the vector
(27) $\phi_{i}=\big{(}W^{(\ell)}_{1}\mathinner{|\phi_{v}\rangle},W^{(\ell)}_{2}\mathinner{|\phi_{v}\rangle},\dots,W^{(\ell)}_{m}\mathinner{|\phi_{v}\rangle}\big{)}^{\dagger}\,.$
It is then clear that $M^{(l)}=\sum_{i}p_{i}\phi^{\dagger}_{i}\phi_{i}\succeq 0$. Likewise a matrix defined by
(28) $G^{(\ell-q_{k})}_{ij}=\operatorname{tr}({W_{i}^{\dagger}}^{(l)}g_{k}(X_{1},\dots,X_{m})W_{j}^{(l)}\varrho)\,,$
with $q_{k}:=\lceil\frac{\text{deg}(g_{k})}{2}\rceil$ is positive semidefinite when $g_{k}(X_{1},\dots,X_{m})$ is positive semidefinite. This can be seen from the fact that a matrix $g_{k}\succeq 0$ is decomposable as a product $g_{k}=b_{k}b_{k}^{\dagger}$ of a matrix $b_{k}$.
Note that $G^{(\ell-q_{k})}$ are sub-matrices of $M^{(\ell)}$.
The optimization problem from Eq. (25) is then approximated by the following semidefinite programming hierarchy indexed by $\ell\in\mathds{N}$ [33, 54, 53],
(29) $\alpha_{\ell}=\max_{M^{(\ell)}}$ $\langle\,f(x_{1},\dots,x_{m})\,\rangle\,,$
subject to $M^{(\ell)}_{00}=1\,,$
$M^{(\ell)}\succeq 0\,,$
$G^{(\ell-q_{k})}\succeq 0\,,$
$M^{(\ell)}_{ij}=M^{(\ell)}_{\alpha\beta}\quad\quad\text{if}\quad{\langle w_{i}^{*}w_{j}\rangle}={\langle w_{\alpha}^{*}w_{\beta}\rangle}\,,$
$G^{(\ell-q_{k})}_{ij}=M^{(\ell)}_{\alpha\beta}\quad\text{if}\quad{\langle w_{i}^{*}g_{k}w_{j}\rangle}={\langle w_{\alpha}^{*}w_{\beta}\rangle}\,,$
for $1\leq k\leq r$ and with $\ell$ greater than or equal to the degree of the polynomial $f$.
Generally $\alpha_{\ell}\neq\alpha^{*}$ since the constraints from Eq. (29) do not imply that $\alpha_{\ell}$ comes from a state $\varrho$ and operators $X_{i}$ satisfying the constraints $g_{i}\succeq 0$. However, when the Archimedeanity condition is met, i.e., there exists a $N\geq 0$ such that $N-\sum^{m}_{i=0}x^{*}_{i}x_{i}\geq 0$, the relaxation converges to the optimum $\alpha^{*}$ of Eq. (25[33, Proposition 6.7],
(30) $\lim_{\ell\rightarrow\infty}\alpha_{\ell}=\alpha^{*}\,.$
A variant of the Gelfand-Naimark-Segal (GNS) construction then allows us to construct a direct sum of states $\varrho$ and operators $\{X_{i}\}$ on a Hilbert space $\mathcal{H}$ from the sequence $\{M^{(\ell)}\}_{\ell=0}^{\infty}$ such that each block satisfies the constraints of the optimization problem [33, 68].
3.7. Quasi-Clifford algebra
Let $F$ be a field of characteristic not equal to $2$ and $m$ a positive integer. A quasi-Clifford (QC) algebra is defined as an associative algebra over $F$ containing the identity $1$ on $m$ generators $\alpha_{1},\dots,\alpha_{m}$, such that $\alpha_{i}^{2}=k_{i}$ and $\alpha_{j}\alpha_{i}=(-1)^{\delta_{ij}}\alpha_{i}\alpha_{j}$ with $k_{i}\in F$. Denote by $\mathbb{C}_{b}$ a QC algebra on one generator $\beta$ with $\beta^{2}=b$ and by $\mathbb{Q}_{cd}$ a QC algebra on two generators $\gamma$, $\delta$ with $\gamma^{2}=c$, $\delta^{2}=d$, and $\gamma\delta=-\delta\gamma$.
Theorem 3.
[21, Theorem 2.7] A QC algebra $\mathscr{C}$ with $m$ generators is isomorphic to
where $m=r+2s$ with $r$, $s>0$ and $b_{i}$, $c_{j}$, $d_{k}$ are plus or minus products of some $k_{i}$.
When $k_{i}=1$ for all $i$ and $F$ is the field of complex numbers, then the QC algebra $\mathscr{C}$ contains only ±$\mathbb{C}_{\pm 1}$ and ±$\mathbb{Q}_{\pm 1\pm 1}$ as tensor factors. Ref. [21, Eq. (3.1)] lists the irreducible representation of the generators of these algebras
(32) $\begin{matrix}[2]\text{Field}&\text{Generators}\\ \hline\cr\mathbb{C}_{1}&\beta\rightarrow 1\quad\text{or}\quad\beta\rightarrow-1\\ \mathbb{C}_{-1}&\beta\rightarrow i\quad\text{or}\quad\beta\rightarrow-i\\ \mathbb{Q}_{\pm 1,1}&\gamma\rightarrow\begin{pmatrix}[1]0&\pm 1\\ 1&0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}[1]1&0\\ 0&-1\end{pmatrix}\\ \mathbb{Q}_{1,-1}&\gamma\rightarrow\begin{pmatrix}[1]0&\phantom{-}1\\ 1&\phantom{-}0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}[1]0&-1\\ 1&0\end{pmatrix}\\ \mathbb{Q}_{-1,-1}&\gamma\rightarrow\begin{pmatrix}[1]0&-1\\ 1&0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}[1]i&0\\ 0&-i\end{pmatrix}\\ \end{matrix}$
Furthermore, Ref. [21, Eq. (2.4)] also shows that ±$\mathbb{C}_{\pm 1}\cong 2F$ and ±$\mathbb{Q}_{\pm 1\pm 1}\cong F_{2}$, where $2F=F\oplus F$ and $F_{2}$ is the algebra of complex $2\times 2$ matrices. This allows us to state the following corollary.
Corollary 4.
Let $\{\alpha_{1},\dots,\alpha_{m}\}$ be a set of $m$ operators from a Hilbert space over $\mathbb{C}$ with $\alpha^{2}_{i}=1$ and $\alpha_{j}\alpha_{i}=(-1)^{\delta_{ij}}\alpha_{i}\alpha_{j}$ with $\delta_{ij}\in\{0,1\}$. Then, these operators are isomorphic, up to phases, to a direct sum of $2^{r}$ $s$-qubit error basis with $m=r+2s$.
Proof.
The set of operators $\{\alpha_{1},\dots,\alpha_{m}\}$ defines a QC algebra $\mathscr{C}$ with $k_{i}=1$ for all $i$. Therefore, according to Theorem 3,
with $b_{i}$, $c_{j}$, $d_{k}\in\{-1,1\}$. Using the fact that ±$\mathbb{C}_{\pm 1}\cong 2F$ and ±$\mathbb{Q}_{\pm 1\pm 1}\cong F_{2}$, one has
Here we used that $\otimes$ distributes over $\oplus$ and $F_{m}\otimes F_{n}\cong F_{nm}$. Note that $F_{2^{r}}$ is the tensor product of $2^{s}$ complex matrices of size $2\times 2$. Therefore, these operators $\{\alpha_{1},\dots,\alpha_{m}\}$ are isomorphic to a direct sum of $s$-qubit operators. According to Eq. (32), these operators are given by a tensor product of Pauli matrices [see Eq. (7)], up to phases. This ends the proof. ∎
The direct sum in Eq. (34) comes from the fact that given a set of operators with some commutation/anti-commutation relations and squaring to $\mathds{1}$, the mapping to Pauli operators is not unique. For example, given $x_{1},x_{2},x_{3}$ with $x_{i}x_{j}=-x_{j}x_{i}$ and $x^{2}_{i}=\mathds{1}$, the elements $x_{1}$, $x_{2}$, $x_{3}$ can be mapped to the Pauli matrices $X$, $Z$, $Y$ but also to $X$, $Z$, $-Y$, respectively. Different blocks from the direct sum give thus different mappings to Pauli operators.
@@ 3.8. Semidefinite programming
where $A_{1},\dots,A_{n},C\in S^{k}$ and $b_{i}\in\mathds{R}$. Eq. (3.8) is called the primal program and $\operatorname{tr}(CX)$ is its objective function. To each primal program is associated a dual program which reads
@@
A variable $X$ (or set of $y_{i}$) satisfying the constraints of the primal (or dual) problem is said to be primal (or dual) feasible. Any pair of feasible $X$ and $\{y_{i}\}$ satisfy the weak duality:
@@
An optimization problem is converted to a feasibility problem by setting $C=0$. The weak duality can then be used to prove primal infeasibility: if there exists a dual solution $\{y_{i}\}$ with $\sum^{p}_{i=1}y_{i}b_{i}\geq 0$, then Eq. (37) is violated and the primal problem is infeasible.
@@
where $\operatorname{Re}(X)$ and the $\operatorname{Im}(X)$ are the real and the imaginary part of $X$, respectively. Then $X\succeq 0$ if and only if $\tilde{X}\succeq 0$

### v2

@@ 3.5. Lovász theta number of a graph
In other words, $\vartheta^{\prime}(G)=\eta_{C}$.
@@ 3.6. Semidefinite programming
where $A_{1},\dots,A_{p},C\in S^{k}$ and $b_{i}\in\mathds{R}$. Eq. (3.6) is called the primal program and $\operatorname{tr}(CX)$ is its objective function. To each primal program is associated a dual program which reads
@@
A variable $X$ (or set of $y_{i}$) satisfying the constraints of the primal (or dual) problem is said to be primal (or dual) feasible. Any pair of feasible $X$ and $\{y_{i}\}_{i=1}^{p}$ satisfy weak duality:
@@
An optimization problem is converted to a feasibility problem by setting $C=0$. Weak duality can then be used to prove primal infeasibility: if there exists a dual feasible solution $\{y_{i}\}_{i=1}^{p}$ with $\sum^{p}_{i=1}y_{i}b_{i}>0$, then Eq. (27) is violated and the primal problem is infeasible.
@@
where $\operatorname{Re}(X)$ and the $\operatorname{Im}(X)$ are the real and the imaginary part of $X$, respectively. Then $X\succeq 0$ if and only if $\tilde{X}\succeq 0$.
3.7. State polynomials
Given a state $\varrho$ and a set of observables $\{X_{i}\}^{m}_{i=1}$, denote by ${\langle X_{i}\rangle}_{\varrho}=\operatorname{tr}(X_{i}\varrho)$ their expectation values 777 Technically, ${\langle X_{i}\rangle}_{\varrho}=\operatorname{tr}(X_{i}\varrho)$ only holds for bounded linear trace-class operators $\varrho$ on a separable Hilbert space. In general, ${\langle X_{i}\rangle}_{\varrho}=\varrho(X_{i})\in\mathds{C}$.. Noncommutative polynomials over $\mathds{C}$ in the products of operators and their expectations are termed state polynomials. An example is ${\langle X_{1}^{2}\rangle}_{\varrho}X_{2}X_{1}+{\langle X_{1}\rangle}_{\varrho}{\langle X_{2}X_{1}X_{2}\rangle}_{\varrho}$ where the second term is understood as $\mathds{1}{\langle X_{1}\rangle}_{\varrho}{\langle X_{2}X_{2}X_{2}\rangle}_{\varrho}$. Polynomials containing “bare” operators (e.g., ${\langle X_{1}^{2}\rangle}_{\varrho}X_{2}X_{1}$) are non-commutative state polynomials, whereas those only containing expectation values (e.g., ${\langle X_{1}\rangle}_{\varrho}{\langle X_{2}X_{1}X_{2}\rangle}_{\varrho}$) are commutative or pure. How can one optimize state polynomials over the set of all states and operators satisfying a given set of constraints?
Let $f$ and $\{g_{j}\}_{j=1}^{r}$ be state polynomials. Consider optimizing $f$ under positive semidefinite constraints $g_{k}\succeq 0$ over the set of quantum states $\varrho$ and operators $X_{i}$ acting on a separable Hilbert space $\mathcal{H}$,
(29) $\zeta^{*}=\inf_{\mathcal{H},\{X_{i}\},\varrho}$ $\big{\langle}\,f(X_{1},\dots,X_{m})\,\big{\rangle}_{\varrho}$
subject to $g_{k}(X_{1},\dots,X_{m})\succeq 0\quad\text{for}\quad k=1,\dots,r\,.$
Note that Eq. (29) requires an optimization over the Hilbert space $\mathcal{H}$ and all operators and quantum states with support on $\mathcal{H}$ satisfying the positive semidefinite constraints. The framework of state polynomial optimization [34] (also known as scalar extension [57, 58]) provides a hierarchy of semidefinite programming approximations to Eq. (29).
To see how this works, consider a set of (non-commutative) letters $\{x_{i}\}^{m}_{i=1}$ and form words $w=x_{a_{1}}\dots x_{a_{\xi}}$. Denote the empty word as $1$, satisfying $1w=w1=w$ for all words $w$. The degree of a word is defined as the number of letters in it. The involution of a word is given by $(x_{1}\cdots x_{m})^{*}=(x^{*}_{m}\cdots x^{*}_{1})$. We associate to each word $w$ a scalar with the symbol $\langle w\rangle$}. We refer to $\langle w\rangle$} as the “expectation value” of a word. Expectations satisfy
$w{\langle v\rangle}$ $={\langle v\rangle}w\,,\quad\quad\quad\quad{\langle w{\langle v\rangle}\rangle}={\langle w\rangle}{\langle v\rangle}$
(30) ${\langle w\rangle}{\langle v\rangle}$ $={\langle v\rangle}{\langle w\rangle}\,,\quad\quad\quad({\langle v\rangle}w)^{*}={\langle v^{*}\rangle}w^{*}\,$,
for all words $v,w$  [^8] Note the difference to trace polynomials, which additionally satisfy $\operatorname{tr}(vw)=\operatorname{tr}(wv)$. However, state polynomials generally do not satisfy ${\langle vw\rangle}={\langle wv\rangle}$.. In general, $x_{i}^{*}$ are considered separate variables from the $x_{i}$. They are related to each other by their real and imaginary parts, so that ${\langle x_{i}^{*}\rangle}^{\text{Re}}={\langle x_{i}\rangle}^{\text{Re}}$ and ${\langle x_{i}^{*}\rangle}^{\text{Im}}=-{\langle x_{i}\rangle}^{\text{Im}}$. This allows one to define ${\langle v^{*}\rangle}=\mkern 1.5mu\overline{\mkern-1.5mu{\langle v\rangle}\mkern-1.5mu}\mkern 1.5mu$. In case of Hermitian variables one identifies $x_{i}=x_{i}^{*}$.
Products of words and their expectations are known as (letter) state monomials, e.g., $w_{i_{1}}{\langle w_{i_{2}}\rangle}\dots\langle w_{i_{m}}\rangle$. The degree of a state monomial is the sum over the degree of all words that compose it, including the words within expectations. State polynomials are formed by a linear combination of monomials over $\mathds{C}$, where the expectation $\langle\,\cdot\,\rangle$ is extended linearly. The degree of a state polynomial is the largest degree of the monomials that it is composed of. Let $\mathds{C}\langle x\rangle$ be the set of non-commutative polynomials over $\mathds{C}$, and $\mathds{C}\langle x\rangle_{k}$ its subset of degree at most $k$. Let $\bf S$ be the set of non-commutative state polynomials over $\mathds{C}$, and $S=\{\langle p\rangle\,|\,p\in\bf{S}\}$ be the set of (commutative/pure) state polynomials.
Naturally, (letter) state polynomials behave in the same manner as polynomials in operators and their expectations, i.e., operator state polynomials. A word $w=x_{a_{1}}\cdots x_{a_{\xi}}$ and its expectation value $\langle w\rangle$} can be evaluated on a state $\varrho$ and operators $X_{a_{1}},\dots,X_{a_{\xi}}$ so that $W=X_{a_{1}}\cdots X_{a_{\xi}}$ and ${\langle W\rangle}_{\varrho}={\langle X_{a_{1}}\cdots X_{a_{\xi}}\rangle}_{\varrho}$. Then, the empty word $1$ corresponds to the identity $\mathds{1}$ and ${\langle 1\rangle}=1$. A monomial $(x_{a_{1}}\cdots x_{a_{\xi}}){\langle x_{b_{1}}\cdots x_{b_{\mu}}\rangle}\cdots{\langle x_{\ell_{1}}\cdots x_{\ell_{\nu}}\rangle}$ is evaluated as $(X_{a_{1}}\cdots X_{a_{\xi}}){\langle X_{b_{1}}\cdots X_{b_{\mu}}\rangle}_{\varrho}\cdots{\langle X_{\ell_{1}}\cdots X_{\ell_{\nu}}\rangle}_{\varrho}$. By linear extension any state polynomial can be evaluated.
In what follows we use the term state polynomials loosely to denote either type, that is letter or operator state polynomials. The important difference between the two types is that when assigning a set of values to the $\langle w\rangle$}, there does not necessarily exists a compatible quantum state and operators. Thus with some abuse of notation, given a set of $\langle w\rangle$$\in\mathds{C}$, it can be that there does not exist a $\varrho$ and a set of operators $W$ such that ${\langle w\rangle}={\langle W\rangle}_{\varrho}$.
3.8. State polynomial optimization
Let $\varrho$ be a state and $S_{0},\dots,S_{u}$ be the set of all operator state monomials of degree at most $\ell\in\mathds{N}$ in operators $X_{1},\dots,X_{m}$. We set $S_{0}=\mathds{1}$ for simplicity. One can then form a moment matrix with entries
(31) $M^{(\ell)}_{ij}=\operatorname{tr}\big{(}{S_{i}^{\dagger}}S_{j}\varrho\big{)}\quad\in\mathds{C}^{(u+1)\times(u+1)}\,.$
Then $M_{00}=1$ and $M^{(\ell)}_{ij}=M^{(\ell)}_{rs}$ if $\langle S_{i}^{\dagger}S_{j}\rangle=\langle S_{r}^{\dagger}S_{t}\rangle$. The moment matrix $M^{(\ell)}$ is positive semidefinite. This can be seen from the spectral decomposition $\varrho=\sum_{v}\lambda_{v}|\phi_{v}\rangle\langle\phi_{v}|$ with $\lambda_{v}\geq 0$; then for each $\mathinner{|\phi_{v}\rangle}$ define the vector
(32) $\phi_{v}=\big{(}S_{0}\mathinner{|\phi_{v}\rangle},S_{1}\mathinner{|\phi_{v}\rangle},\dots,S_{u}\mathinner{|\phi_{v}\rangle}\big{)}^{\dagger}\,.$
One sees that $M^{(\ell)}=\sum_{v}\lambda_{v}\phi_{v}\phi_{v}^{\dagger}\succeq 0$.
Likewise, given a state polynomial inequality $g_{k}(X_{1},\dots,X_{m})\succeq 0$, we can form a moment matrix using all state monomials of up to degree $\ell-q_{k}$ where $q_{k}:=\lceil\frac{\text{deg}(g_{k})}{2}\rceil$  [^9] Choose $\ell$ large enough so that $\ell-q_{k}\geq 0$.,
(33) $G^{(\ell-q_{k})}_{ij}=\operatorname{tr}({S_{i}^{\dagger}}g_{k}(X_{1},\dots,X_{m})S_{j}\varrho)\,$,
Also $G^{(\ell-q_{k})}_{ij}$ is positive semidefinite, which follows from the fact that for every feasible point, $g_{k}\succeq 0$ can be factorized as a product $g_{k}=b_{k}b_{k}^{\dagger}$ of a matrix $b_{k}$. Note that the entries of $G^{(\ell-q_{k})}$ are linear combinations of the entries of $M^{(\ell)}$.
The idea of state polynomial optimization is to optimize over a semidefinite program where $M^{(\ell)}$ and the matrices $G^{(\ell-q_{k})}$, $k=1,\dots,r$ are positive semidefinite matrix variables, such that they satisfy all linear constraints arising from relations among the entries of $M^{(\ell)}$ and the $G^{\ell-q_{k}}$. This provides a relaxation for the optimization of a state polynomial under a set of state polynomial constraints.
To formulate the optimization problem, one uses letter state polynomials. For $s,t$ state polynomials we write the entries of the state moment matrix as $M^{(\ell)}_{st}={\langle s^{*}t\rangle}$. To simplify the notation, define $M^{(\ell)}({\langle s^{*}t\rangle}):=M^{(\ell)}_{st}$.
The optimization problem from Eq. (29) is then approximated by the following semidefinite programming hierarchy [34, 58, 57] indexed by $\ell\in\mathds{N}$,
(34) $\zeta_{\ell}\quad=\quad\inf_{M^{(\ell)}}$ $\langle\,f(x_{1},\dots,x_{m})\,\rangle\,$,
subject to $M^{(\ell)}(1)=1\,$,
$M^{(\ell)}\succeq 0\,$,
$G^{(\ell-q_{k})}\succeq 0\quad\quad\quad\quad\quad\quad\,\,\text{for}\quad 1\leq k\leq r\,$,
$M^{(\ell)}_{st}=M^{(\ell)}_{\alpha\beta}\quad\quad\quad\quad\quad\quad\text{if}\quad\quad{\langle s^{*}t\rangle}={\langle\alpha^{*}\beta\rangle}\,$,
$G^{(\ell-q_{k})}_{st}=\sum_{\alpha\beta}c_{\alpha\beta}M^{(\ell)}_{\alpha\beta}\quad\quad\text{if}\quad\quad{\langle s^{*}g_{k}t\rangle}=\sum_{\alpha\beta}c_{\alpha\beta}{\langle\alpha^{*}\beta\rangle}\,$,
where $r$ is the number of constraints, $\ell$ is greater than or equal to the degree of the polynomial $f$, and $c_{\alpha\beta}\in\mathds{C}$.
The program (34) provides a lower bound $\zeta_{\ell}\leq\zeta^{*}$, but in general $\zeta_{\ell}\neq\zeta^{*}$. Let us consider convergence of $\zeta_{\ell}$ to $\zeta^{*}$. A constraint set is balanced, if $C^{*}=C$ and for every non-symmetric element $g_{k}^{*}\neq g_{k}$ we also have $g_{k}\in C$. This allows us to handle equality constraints by requiring both $g_{k}\succeq 0$ and $g_{k}\succeq 0$. Corollary 6.1 in combination with Corollary 6.7 in [34] states that if the constraint set is balanced and if there exists a $N>0$ such that $N-\sum_{i}x_{i}^{*}x_{i}=\sum_{j}p_{j}^{*}p_{j}+\sum_{k}\lambda_{k}g_{k}$ for some $p_{j}\in\mathds{C}\langle x\rangle_{1}$, $g_{k}\in C\cap\mathds{C}\langle x\rangle_{2}$, and $\lambda_{k}§\in\mathds{R}_{\geq 0}$, then the sequence of relaxations in Eq. (34) converges to the optimum $\zeta^{*}$ of Eq. (29),
(35) $\lim_{\ell\rightarrow\infty}\zeta_{\ell}=\zeta^{*}\,.$
Under certain flatness conditions, a variant of the Gelfand-Naimark-Segal (GNS) construction allows one to construct state $\varrho$ and operators $\{X_{i}\}$ on a Hilbert space $\mathcal{H}$ from the sequence $\{M^{(\ell)}\}_{\ell=0}^{\infty}$ satisfying the constraints of the optimization problem [34, Proposition 6.10].
A short remark on imposing constraints: inequality constraints are imposed by building a localizing matrix $G^{(\ell-q_{k})}_{st}\succeq 0$ from each constraint $g_{k}$. Equality constraints $g_{i}=0$ can be imposed in two different ways: either they are enforced by requiring both $g_{k}\succeq 0$ and $g_{k}\succeq 0$, or alternatively, a set of equalities $E=\{e_{i}\,|\,i=1,\dots,m_{e}\}$ can be modeled by replacing $\bf S$ by the quotient ring ${\bf S}/I$, where $I=\left\{\sum_{i}p_{i}g_{i}q_{i}:p_{i},q_{i}\in{\bf S}\right\}$. A simplified construction of this can be done for group algebras [34, Section 7.1].
3.9. Quasi-Clifford algebras
Let $F$ be a commutative field of characteristic not equal to $2$ and $m$ be a positive integer. Let $\{k_{j}\}_{j=1}^{m}\in F$ be non-zero elements and $\chi_{ij}\in\{0,1\}$ for $1\leq i<j\leq m$. A quasi-Clifford (QC) algebra is defined as an associative algebra over $F$ containing the identity formed by $m$ generators $\alpha_{1},\dots,\alpha_{m}$, such that
(36) $\alpha_{i}^{2}=k_{i}\,,\quad\quad\alpha_{j}\alpha_{i}=(-1)^{\chi_{ij}}\alpha_{i}\alpha_{j}\quad\text{for }i<j\,.$
Denote by $\mathbb{C}_{b}$ the QC algebra of one generator $\beta$ with $\beta^{2}=b$ and by $\mathbb{Q}_{cd}$ the QC algebra of two generators $\gamma$, $\delta$ with $\gamma^{2}=c$, $\delta^{2}=d$, and $\gamma\delta=-\delta\gamma$.
Theorem 3.
[21, Theorem 2.7] A QC algebra $\mathscr{C}$ with $m$ independent generators is isomorphic to
where $m=r+2s$ with $r$, $s\geq 0$ and $b_{i}$, $c_{i}$, $d_{i}$ are plus or minus products of some $k_{j}$’s.
Restricting to $\alpha_{i}^{2}=k_{i}=1$ and $F$ to $\mathds{C}$, then $\mathscr{C}$ contains only $\mathbb{C}_{\pm 1}$ and $\mathbb{Q}_{\pm 1\pm 1}$ as tensor factors. Ref. [21, Eq. (3.1)] lists the irreducible representations of the generators.
(38) $\begin{array}[]{ll}\text{Field}&\text{Generators}\\[2.84544pt] \hline\cr\\[-11.38092pt] \mathbb{C}_{1}&\beta\rightarrow 1\quad\text{or}\quad\beta\rightarrow-1\\[5.69046pt] \mathbb{C}_{-1}&\beta\rightarrow i\quad\text{or}\quad\beta\rightarrow-i\\[5.69046pt] \mathbb{Q}_{\pm 1,1}&\gamma\rightarrow\begin{pmatrix}0&\pm 1\\ 1&0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\\[14.22636pt] \mathbb{Q}_{1,-1}&\gamma\rightarrow\begin{pmatrix}0&\phantom{-}1\\ 1&\phantom{-}0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}[1]0&-1\\ 1&0\end{pmatrix}\\[14.22636pt] \mathbb{Q}_{-1,-1}&\gamma\rightarrow\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\quad\delta\rightarrow\begin{pmatrix}i&0\\ 0&-i\end{pmatrix}\end{array}$
This shows that $\mathbb{C}_{\pm 1}\cong 2F$ and $\mathbb{Q}_{\pm 1,\pm 1}\cong F_{2}$, where $2F=F\oplus F$ and $F_{2}$ is the algebra of complex $2\times 2$ matrices. Thus, Eq. (37) can be written as a direct sum
Here we use that $\otimes$ distributes over $\oplus$ and $F_{m}\otimes F_{n}\cong F_{nm}$. As a consequence, any quasi-Clifford algebra over $\mathds{C}$ with $k_{i}=1$
is isomorphic to a direct sum of $2^{r}$ algebras, each formed (up to phases) by $s$-qubit Pauli operators, where $m=r+2s$.
Remark 4.
The direct sum in Eq. (39) comes from the fact that given a set of operators with some commutation/anti-commutation relations and squaring to $\mathds{1}$, the mapping to Pauli operators is not unique. For example, given $x_{1},x_{2},x_{3}$ with $x_{i}x_{j}=-x_{j}x_{i}$ and $x_{i}^{2}=\mathds{1}$, the elements $x_{1}$, $x_{2}$, $x_{3}$ can be mapped to the Pauli matrices $X$, $Z$, $Y$, respectively, but also to $X$, $Z$, Y-Y, respectively. Different blocks in the direct sum correspond to different inequalivalent mappings to Pauli operators.